News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 13 June 2003 8.00am EDT

During the last week Chandra completed the observing schedule as planned.

A real-time procedure was completed on June 12 to perform a warm re-start of the noisy IRU-1 gyro-1. On re-start, the IRU-1 gyro-1 current reset from its recent level of ~140 mA to 96 mA, a value comparable to its pre-launch level. The reset of the current is very promising and its future trend will be monitoring carefully. IRU-2 continues in a powered on state in parallel with operations and on June 9 a series of calibration maneuvers were executed from the daily loads.

A momentum dump was performed on June 12 prior to the gyro-1 warm start and was completed with nominal thruster performance. The unload was performed at a relatively low starting temperature.

A third real-time procedure was run on June 12 to lower the ACA CCD temperature from -12 to -13 deg C to mitigate the increasing number of warm pixels. This is the third of five planned 1 degree step reductions.

Of note last week was the conduct of the Chandra User's Committee meeting on Jun 12-13.

The observing schedule for next week is shown below and includes two observations of NGC 526A coordinated with XMM.

------------------------------------------
Eta Carinae            ACIS-S/HETG  Jun 16
2MASS042032-204759     ACIS-S       Jun 17
BR 0331-1622           ACIS-S
SDSS J022756+005733.1  ACIS-S
Radiation Belts
1E 1207.4-5209         ACIS-S/LETG  Jun 18
X1702-429              HRC-I        Jun 19
X1715-321              HRC-I
SGR A*                 ACIS-I
XTE J1814-338          ACIS-S/HETG
Radiation Belts                     Jun 20
SDSS J022841+005208.6  ACIS-S
SDSS J2215-0045        ACIS-S
NGC 526A               ACIS-S/HETG
KES 27                 ACIS-I       Jun 21
NGC 526A               ACIS-S/HETG
SN 2003BG              ACIS-S       Jun 22

------------------------------------------

All spacecraft subsystems, with the exception of the MUPS thruster A-1 under investigation, continued to operate nominally.

Return to Status Reports